you must prepare a 100mL solution of hydrochloric acid with a concentration of 0.500mol/L. You are given the appropriate glassware and a bottle of 2.50mol/L stock hydrochloric acid. What volume of stock solution must you use to prepare the solution?
1
Expert's answer
2020-06-08T15:27:56-0400

V1 - ?; С1 = 2.50 mol/L, C2 = 0.500 mol/L, V2 = 0.100 L


C1 = n/V1 ; С2 = n/V2 "\implies" C1V1 = C2V2


V1 = C2V2/C1 = 0.500*0.100/2.50 = 0.0200 L = 20.0 mL